Computing the spectrum of the one-dimensional p-Laplacian

B Malcolm Brown Cardiff University

The talk will focus on some recent work to compute the eigenvalues and Fucik spectrum of the one-dimensional p-Laplacian. This is joint work with Wolfgang Reichel of Basel.